# On a class of thin obstacle-type problems for the bi-Laplacian operator

### v2

@@ 4. Regularity of solutions
$v\in L^{\beta_{0}2^{**}_{b}/2}(B_{\rho_{0}}^{+};y^{b})\cap L^{\beta_{0}2^{*}_{b}/2}(B_{\rho_{0}}^{\prime}),$ (4.22)
@@
    If $b\geq 0$ then $(u,v)\in C^{2,\alpha}(B_{r}^{+})\times C^{0,\alpha}(B_{r}^{+})$ for all $\alpha\in(0,1-b)$.
@@
    If $b<0$ then $(u,v)\in C^{3,\alpha}(B_{r}^{+})\times C^{1,\alpha}(B_{r}^{+})$ for all $\alpha\in(0,-b)$.
@@
Fix $r<r_{2}<r_{1}<1$. By ˜4.5 (see in particular (4.4)), together with (H.2), we obtain that $f(\cdot,u(\cdot))\in L^{q}(B_{r_{1}}^{\prime})$ for $q\in(n/(1-b),n/(1-b)+1]$ if $b\geq 0$ and $q\in(-n/b,-n/b+1]$ if $b<0$. Thus, we are in a position to apply ˜2.8 and conclude that $v\in C^{0,\alpha}(B_{r_{2}}^{+})$ for all $\alpha\in(0,1-b)$ whenever $b\geq 0$, and $v\in C^{1,\alpha}(B_{r_{2}}^{+})$ for all $\alpha\in(0,-b)$ if $b<0$. In turn, by ˜2.10, we obtain that $u\in C^{2,\alpha}(B_{r}^{+})$ with $\alpha\in(0,1-b)$ for $b\geq 0$, and $u\in C^{3,\alpha}(B_{r}^{+})$ with $\alpha\in(0,-b)$ when $b<0$. ∎
@@
    Assume that $b\geq 0$. Then for all $0<r<R<1$ and all $\alpha\in(0,1-b)$ there exists a constant $C=C(n,b,R,r,\alpha)$ such that
    $\|\nabla_{x}^{\prime}v\|_{C^{0,\alpha}(B_{r}^{+})}\leq C\left(\|v\|_{H^{1}(B_{R}^{+};y^{b})}+1+\|\nabla_{x}^{\prime}u\|_{L^{q}(B_{R}^{\prime})}\right).$
@@
    Assume that $b<0$. Then for all $0<r<R<1$ and all $\alpha\in(0,-b)$ there exists a constant $C=C(n,b,R,r,\alpha)$ such that
    $\|\nabla_{x}^{\prime}v\|_{C^{1,\alpha}(B_{r}^{+})}\leq C\left(\|v\|_{H^{1}(B_{R}^{+};y^{b})}+1+\|\nabla_{x}^{\prime}u\|_{L^{q}(B_{R}^{\prime})}\right).$
Moreover, there exists $\beta\in(0,1)$ such that $\partial_{y}^{b}v\in C^{0,\beta}(B_{r}^{+})$.
@@
$\|f^{\xi}\|_{L^{q}(B_{r_{1}}^{\prime})}^{q}\leq C\left(1+\int_{B_{r_{1}}^{\prime}}\left|\frac{u(x+\xi,0)-u(x,0)}{|\xi|}\right|^{q}\,dx\right)\leq C\left(1+\|\nabla_{x}^{\prime}u\|_{L^{q}(B_{R}^{\prime})}^{q}\right).$ (4.23)
In particular, we are in a position to apply ˜2.8 and conclude that if $b\geq 0$ then $v^{\xi}\in C^{0,\alpha}(B_{r}^{+})$ with
$\|v^{\xi}\|_{C^{0,\alpha}(B_{r}^{+})}$ $\leq C\left(\|v^{\xi}\|_{L^{2}(B_{r_{1}}^{+};y^{b})}+\|f^{\xi}\|_{L^{q}(B_{r_{1}}^{\prime})}\right)$
$\leq C\left(\|v\|_{H^{1}(B_{R}^{+};y^{b})}+1+\|\nabla_{x}^{\prime}u\|_{L^{q}(B_{R}^{\prime})}\right)$ (4.24)
for all $\alpha\in(0,1-b)$, where in the last step we have used (4.23) and a classical estimate for the difference quotient of Sobolev functions. Observe that the inequality obtained in (4.24) is uniform with respect to $\xi$. Therefore, the Ascoli–Arzelá Theorem can then be used to show that $\nabla_{x}^{\prime}v\in C^{0,\alpha}(B_{r}^{+})$ for all $\alpha\in(0,1-b)$. A similar argument yields the desired result when $b<0$.
Finally, since $x\mapsto f(x,u(x,0))$ is a bounded Lipschitz function, then it belongs to $C^{0,\sigma}(B_{R}^{\prime})$ for all $\sigma\in(0,1)$. In particular, ˜2.9 implies that $\partial_{y}^{b}v\in C^{0,\beta}(B_{r}^{+})$ for some $\beta\in(0,1)$. This concludes the proof. ∎
@@
Under the assumptions of ˜4.7, assume that $b=0$ and let $r<1$. Then, there exists $\beta\in(0,1)$ such that $(u,v)\in C^{3,\beta}(B_{r}^{+})\times C^{1,\beta}(B_{r}^{+})$.
@@
Fix $r<R<1$ and observe that by ˜4.7, when $b=0$, we have that $v\in C^{1,\beta}(B_{R}^{+})$ for some $\beta\in(0,1)$. The desired conclusion then follows from classical regularity theory (see also ˜2.10 $(i)$). ∎
@@
$\left\{\begin{array}[]{rl}u^{(iv)}=&0,\\ u^{\prime\prime\prime}(0)=&-6=\lambda_{-}u^{-}(0)-\lambda_{+}u^{+}(0),\\ u^{\prime}(0)=&0.\end{array}\right$.
Moreover $u>0$ in $(0,1)$, but $u^{\prime\prime}(x)=-6x<0$, thus showing that $u$ is superharmonic.
@@ 5. Almgren’s frequency formula
$D_{0}(r,u,v)\coloneqq\int_{B_{r}^{+}}y^{b}(|\nabla u|^{2}+|\nabla v|^{2})\,dX$ (5.2)
@@
$D(r,u,v)\coloneqq\int_{B_{r}^{+}}y^{b}(|\nabla u|^{2}+|\nabla v|^{2}+u\Delta_{b}u)\,dX+\int_{B_{r}^{\prime}}(\partial_{y}^{b}v)v\,dx.$ (5.4)
@@
$\int_{\partial B_{r}^{\prime}}huv\,d\mathcal{H}^{n-1}$ (5.5)
@@
$P(r,u,v)\coloneqq\int_{B_{r}^{+}}y^{b}u\Delta_{b}u\,dX+\int_{B_{r}^{\prime}}(\partial_{y}^{b}v)v\,dx$ (5.6)
@@
$\left|\int_{B_{r}^{+}}y^{b}uv\,dX\right|\leq\frac{1}{2}\int_{B_{r}^{+}}y^{b}(u^{2}+v^{2})\,dX\leq Cr(rD_{0}(r)+H(r)).$ (5.7)
@@
$\left|\int_{B_{r}^{\prime}}f(x,u)v\,dx\right|\leq C\int_{B_{r}^{\prime}}|u|^{p-1}|v|\,dx\leq Cr^{-b}(rD_{0}(r)+H(r)),$ (5.8)
@@
$\int_{B_{R}^{+}}y^{b}\left|u\nabla u\cdot\frac{X}{|X|}+v\nabla v\cdot\frac{X}{|X|}\right|\,dX=\int_{0}^{R}\int_{S_{r}^{+}}y^{b}|uu_{\nu}+vv_{\nu}|\,d\mathcal{H}^{n}\,dr.$ (5.9)
@@
Arguing by contradiction, assume that $u$ and $v$ are not both identically equal to zero in $B_{r_{0}}^{+}$ and that there exists $r<r_{0}$ such that $H(r)=0$. Additionally, in this step we assume that $r$ is a Lebesgue point of $d$. Let $\delta>0$ be such that $r+\delta<r_{0}$. Since by ˜5.3 we have that $D(\rho)=d(\rho)$ for $\mathcal{L}^{1}$-a.e. $\rho\in(r,r+\delta)$ (see also (5.10)), it follows from ˜5.2 that
@@
$\fint_{r}^{r+\delta}(1-C\rho^{1-b})D_{0}(\rho)\,d\rho\leq\fint_{r}^{r+\delta}d(\rho)\,d\rho+\fint_{r}^{r+\delta}C\rho^{-b}H(\rho)\,d\rho.$ (5.11)
@@
Step 2: To complete the proof, it is enough to show that every $r<r_{0}$ is a Lebesgue point for $d$. To this end, observe that by ˜4.7, for every $b\in(-1,1)$ there exists $\alpha>0$ such that $uu_{\nu}+vv_{\nu}\in C^{0,\alpha}(B_{r_{0}}^{+})$; in order to simplify the notation, for the remainder of the proof we use $w$ to denote this function. Observe that, by a change of variables,
@@
This readily implies that $d\in C(0,r_{0})$ and the proof is complete. ∎
@@
$D_{0}^{\prime}(r)=\frac{n+b-1}{r}D_{0}(r)+2\int_{S_{r}^{+}}y^{b}(u_{\nu}^{2}+v_{\nu}^{2})\,d\mathcal{H}^{n}\\ -\frac{2}{r}\int_{B_{r}^{+}}y^{b}(X\cdot\nabla u)v\,dX-\frac{2}{r}\int_{B_{r}^{\prime}}(x\cdot\nabla_{x}^{\prime}v)f(x,u)\,dx$ (5.12)
@@
$D_{0}^{\prime}(r)=\int_{S_{r}^{+}}y^{b}(|\nabla u|^{2}+|\nabla v|^{2})\,d\mathcal{H}^{n}$ (5.14)
@@
$N_{0}^{\prime}(r)=\frac{D_{0}(r)}{H(r)}+\frac{rD_{0}^{\prime}(r)}{H(r)}-\frac{rD_{0}(r)H^{\prime}(r)}{H(r)^{2}}.$ (5.16)
@@
$=-(n+b-1)\frac{D_{0}(r)}{H(r)}+\frac{rD_{0}^{\prime}(r)}{H(r)}-\frac{2rD(r)^{2}}{H(r)^{2}}+\frac{2rP(r)D(r)}{H(r)^{2}}.$ (5.17)
@@
$N_{0}^{\prime}(r)\geq-\frac{2}{H(r)}\int_{B_{r}^{+}}y^{b}(X\cdot\nabla u)v\,dX-\frac{2}{H(r)}\int_{B_{r}^{\prime}}(x\cdot\nabla_{x}^{\prime}v)f(x,u)\,dx+\frac{2rP(r)D(r)}{H(r)^{2}}$ (5.18)
@@
$\frac{2}{H(r)}\int_{B_{r}^{+}}y^{b}(X\cdot\nabla u)v\,dX\geq-C(N_{0}(r)+r^{2}).$ (5.19)
@@
$\int_{B_{r}^{\prime}}(x\cdot\nabla_{x}^{\prime}v)f(x,u)\,dx=\frac{1}{r}\int_{S_{r}^{+}}(y(x\cdot\nabla_{x}^{\prime}v)-|x|^{2}v_{y})f(x,u)\,d\mathcal{H}^{n}\\ +\int_{B_{r}^{+}}[nf(x,u)v_{y}+(x\cdot\nabla_{x}^{\prime}f(x,u))v_{y}+(x\cdot\nabla_{x}^{\prime}u)\partial_{\zeta}f(x,u)v_{y}-(x\cdot\nabla_{x}^{\prime}v)\partial_{\zeta}f(x,u)u_{y}]\,dX.$ (5.20)
@@
$\leq Cr\int_{S_{r}^{+}}y^{b}\left(r|\nabla v|^{2}+\frac{u^{2}}{r}\right)\,d\mathcal{H}^{n},$
@@
$\left|\frac{1}{r}\int_{S_{r}^{+}}(y(x\cdot\nabla_{x}^{\prime}v)-|x|^{2}v_{y})f(x,u)\,d\mathcal{H}^{n}\right|\leq C(r^{2}D_{0}^{\prime}(r)+H(r)).$ (5.21)
@@
$\left|\int_{B_{r}^{+}}nf(x,u)v_{y}\,dX\right|$ $\leq C(rD_{0}(r)+H(r)),$ (5.22)
$\left|\int_{B_{r}^{+}}(x\cdot\nabla_{x}^{\prime}u)\partial_{\zeta}f(x,u)v_{y}-(x\cdot\nabla_{x}^{\prime}v)\partial_{\zeta}f(x,u)u_{y})\,dX\right|$ $\leq CrD_{0}(r),$ (5.23)
$\left|\int_{B_{r}^{+}}(x\cdot\nabla_{x}^{\prime}f(x,u))v_{y}\,dX\right|$ $\leq Cr(D_{0}(r)+rH(r)).$ (5.24)
@@
$\frac{2}{H(r)}\int_{B_{r}^{\prime}}(x\cdot\nabla_{x}^{\prime}v)f(x,u)\,dx\geq-C\left(\frac{r^{2}D_{0}^{\prime}(r)}{H(r)}+1+N_{0}(r)\right)$ (5.25)
@@
$N_{0}^{\prime}(r)\geq-C\left(1+N_{0}(r)+N_{0}(r)^{2}+\frac{r^{2}D_{0}^{\prime}(r)}{H(r)}\right)$ (5.27)
@@
$h(r)\coloneqq\frac{H(r)}{r^{n+b}}=\frac{1}{r^{n+b}}\int_{S_{r}^{+}}y^{b}(u^{2}+v^{2})\,d\mathcal{H}^{n}.$ (5.32)
@@
    where $C$ is a constant that depends on on $\mu$ and on $\|u\|_{L^{\infty}(B_{r_{0}}^{\prime})}^{p-2}$ (when $p>2$). In particular, $h(r)=\mathcal{O}(r^{2\mu})$ as $r\to 0^{+}$.
@@
$\frac{d}{dr}\log(r^{-2\mu}h(r))=\frac{2}{r}\left(N_{0}(r)-\mu+\frac{rP(r)}{H(r)}\right).$ (5.33)
@@
$\geq\frac{2}{r}\left(-\frac{(1-Cr^{1-b})Cr(1+\mu)^{2}}{1+Cr(1+\mu)}-Cr^{1-b}(1+\mu)\right).$ (5.35)

### v3

@@ 4. Regularity of solutions
(4.22) $v\in L^{\beta_{0}2^{**}_{b}/2}(B_{\rho_{0}}^{+};y^{b})\cap L^{\beta_{0}2^{*}_{b}/2}(B_{\rho_{0}}^{\prime}),$
@@
    If $b\geq 0$ then $(u,v)\in C^{2,\alpha}(B_{r}^{+}\cup B^{\prime}_{r})\times C^{0,\alpha}(B_{r}^{+}\cup B^{\prime}_{r})$ for all $\alpha\in(0,1-b)$.
@@
    If $b<0$ then $(u,v)\in C^{3,\alpha}(B_{r}^{+}\cup B^{\prime}_{r})\times C^{1,\alpha}(B_{r}^{+}\cup B^{\prime}_{r})$ for all $\alpha\in(0,-b)$.
@@
Fix $r<r_{2}<r_{1}<1$. By ˜4.5 (see in particular (4.4)), together with (H.2), we obtain that $f(\cdot,u(\cdot))\in L^{q}(B_{r_{1}}^{\prime})$ for $q\in(n/(1-b),n/(1-b)+1]$ if $b\geq 0$ and $q\in(-n/b,-n/b+1]$ if $b<0$. Thus, we are in a position to apply ˜2.8 and conclude that $v\in C^{0,\alpha}(B_{r_{2}}^{+}\cup B^{\prime}_{r_{2}})$ for all $\alpha\in(0,1-b)$ whenever $b\geq 0$, and $v\in C^{1,\alpha}(B_{r_{2}}^{+}\cup B^{\prime}_{r_{2}})$ for all $\alpha\in(0,-b)$ if $b<0$. In turn, by ˜2.10, we obtain that $u\in C^{2,\alpha}(B_{r}^{+}\cup B^{\prime}_{r})$ with $\alpha\in(0,1-b)$ for $b\geq 0$, and $u\in C^{3,\alpha}(B_{r}^{+}\cup B^{\prime}_{r})$ with $\alpha\in(0,-b)$ when $b<0$. ∎
@@
    Assume that $b\geq 0$. Then for all $0<r<R<1$, all $\alpha\in(0,1-b)$, and all $q>n/(1-b-\alpha)$ there exists a constant $C=C(n,b,R,r,q,\alpha)$ such that
    $\|\nabla_{x}^{\prime}v\|_{C^{0,\alpha}({B_{r}^{+}\cup B^{\prime}_{r}})}\leq C\left(\|v\|_{H^{1}(B_{R}^{+};y^{b})}+1+\|\nabla_{x}^{\prime}u\|_{L^{q}(B_{R}^{\prime})}\right).$
@@
    Assume that $b<0$. Then for all $0<r<R<1$, all $\alpha\in(0,-b)$, and all $q>n/(-b-\alpha)$ there exists a constant $C=C(n,b,R,r,q,\alpha)$ such that
    $\|\nabla_{x}^{\prime}v\|_{C^{1,\alpha}(B_{r}^{+}\cup B^{\prime}_{r})}\leq C\left(\|v\|_{H^{1}(B_{R}^{+};y^{b})}+1+\|\nabla_{x}^{\prime}u\|_{L^{q}(B_{R}^{\prime})}\right).$
Moreover, there exists $\beta\in(0,1)$ such that $\partial_{y}^{b}v\in C^{0,\beta}{(B_{r}^{+}\cup B^{\prime}_{r})}$ regardless of the sign of $b$.
@@
(4.23) $\|f^{\xi}\|_{L^{q}(B_{r_{1}}^{\prime})}^{q}\leq C\left(1+\int_{B_{r_{1}}^{\prime}}\left|\frac{u(x+\xi,0)-u(x,0)}{|\xi|}\right|^{q}\,dx\right)\leq C\left(1+\|\nabla_{x}^{\prime}u\|_{L^{q}(B_{R}^{\prime})}^{q}\right).$
In particular, we are in a position to apply ˜2.8 and conclude that if $b\geq 0$ then $v^{\xi}\in C^{0,\alpha}({B_{r}^{+}\cup B^{\prime}_{r}})$ with
$\|v^{\xi}\|_{C^{0,\alpha}({B_{r}^{+}\cup B^{\prime}_{r}})}$ $\leq C\left(\|v^{\xi}\|_{L^{2}(B_{r_{1}}^{+};y^{b})}+\|f^{\xi}\|_{L^{q}(B_{r_{1}}^{\prime})}\right)$
(4.24) $\leq C\left(\|v\|_{H^{1}(B_{R}^{+};y^{b})}+1+\|\nabla_{x}^{\prime}u\|_{L^{q}(B_{R}^{\prime})}\right)$
for all $\alpha\in(0,1-b)$, provided that $q>n/(1-b-\alpha)$; we note that in the last step we have used (4.23) and a classical estimate for the difference quotient of Sobolev functions. Observe that the inequality obtained in (4.24) is uniform with respect to $\xi$. Therefore, the Ascoli–Arzelá Theorem can then be used to show that $\nabla_{x}^{\prime}v\in C^{0,\alpha}({B_{r}^{+}\cup B^{\prime}_{r}})$ for all $\alpha\in(0,1-b)$. A similar argument yields the desired result when $b<0$.
Finally, since $x\mapsto f(x,u(x,0))$ is a bounded Lipschitz function, then it belongs to $C^{0,\sigma}(B_{R}^{\prime})$ for all $\sigma\in(0,1)$. In particular, ˜2.9 implies that $\partial_{y}^{b}v\in C^{0,\beta}({B_{r}^{+}\cup B^{\prime}_{r}})$ for some $\beta\in(0,1)$. This concludes the proof. ∎
@@
Under the assumptions of ˜4.7, assume that $b=0$ and let $r<1$. Then, there exists $\beta\in(0,1)$ such that $(u,v)\in C^{3,\beta}(B_{r}^{+}\cup B^{\prime}_{r})\times C^{1,\beta}(B_{r}^{+}\cup B^{\prime}_{r})$}.
@@
Fix $r<R<1$ and observe that by ˜4.7, when $b=0$, we have that $v\in C^{1,\beta}({B_{R}^{+}\cup B^{\prime}_{R}})$ for some $\beta\in(0,1)$. The desired conclusion then follows from classical regularity theory (see also ˜2.10 $(i)$). ∎
@@
$\left\{\begin{array}[]{rl}u^{(iv)}=&0\qquad\mbox{on }(0,1),\\ u^{\prime\prime\prime}(0)=&-6=\lambda_{-}u^{-}(0)-\lambda_{+}u^{+}(0),\\ u^{\prime}(0)=&0.\end{array}\right$.
Moreover $u>0$ on $(0,1)$, but $u^{\prime\prime}(x)=-6x<0$, thus showing that $u$ is superharmonic.
@@ 5. Almgren’s frequency formula
(5.2) $D_{0}(r,u,v)\coloneqq\int_{B_{r}^{+}}y^{b}(|\nabla u|^{2}+|\nabla v|^{2})\,dX$
@@
(5.4) $D(r,u,v)\coloneqq\int_{B_{r}^{+}}y^{b}(|\nabla u|^{2}+|\nabla v|^{2}+u\Delta_{b}u)\,dX+\int_{B_{r}^{\prime}}(\partial_{y}^{b}v)v\,dx.$
@@
(5.5) $\int_{\partial B_{r}^{\prime}}huv\,d\mathcal{H}^{n-1}$
@@
(5.6) $P(r,u,v)\coloneqq\int_{B_{r}^{+}}y^{b}u\Delta_{b}u\,dX+\int_{B_{r}^{\prime}}(\partial_{y}^{b}v)v\,dx$
@@
(5.7) $\left|\int_{B_{r}^{+}}y^{b}uv\,dX\right|\leq\frac{1}{2}\int_{B_{r}^{+}}y^{b}(u^{2}+v^{2})\,dX\leq Cr(rD_{0}(r)+H(r)).$
@@
(5.8) $\left|\int_{B_{r}^{\prime}}f(x,u)v\,dx\right|\leq C\int_{B_{r}^{\prime}}|u|^{p-1}|v|\,dx\leq Cr^{-b}(rD_{0}(r)+H(r)),$
@@
(5.9) $\int_{B_{R}^{+}}y^{b}\left|u\nabla u\cdot\frac{X}{|X|}+v\nabla v\cdot\frac{X}{|X|}\right|\,dX=\int_{0}^{R}\int_{S_{r}^{+}}y^{b}|uu_{\nu}+vv_{\nu}|\,d\mathcal{H}^{n}\,dr.$
@@
Arguing by contradiction, assume that $u$ and $v$ are not both identically equal to zero in $B_{r_{0}}^{+}$ and that there exists $r<r_{0}$ such that $H(r)=0$. In this step, we assume that $r$ is a Lebesgue point of both $d$ and $H$; this assumption will be removed in the next step by an application of ˜1.2. Let $\delta>0$ be such that $r+\delta<r_{0}$. Since by ˜5.3 we have that $D(\rho)=d(\rho)$ for $\mathcal{L}^{1}$-a.e. $\rho\in(r,r+\delta)$ (see also (5.10)), it follows from ˜5.2 that
@@
(5.11) $\fint_{r}^{r+\delta}(1-C\rho^{1-b})D_{0}(\rho)\,d\rho\leq\fint_{r}^{r+\delta}d(\rho)\,d\rho+\fint_{r}^{r+\delta}C\rho^{-b}H(\rho)\,d\rho.$
@@
Step 2: To complete the proof, it is enough to show that every $r<r_{0}$ is a Lebesgue point for $d$ and $H$. We establish the result for $d$. The argument for $H$ is analogous and involves simpler computations, so we omit the details. To this end, let $\nu=X/|X|$ and observe that by ˜4.7, for $b\leq 0$ there exists $\alpha>0$ such that $uu_{\nu}+vv_{\nu}\in C^{0,\alpha}(B_{r_{0}}^{+}\cup B_{r_{0}}^{\prime})$; in order to simplify the notation, here we use $w$ to denote this function. Note also that, by a change of variables,
@@
This readily implies that $d\in C(0,r_{0})$, and the desired result follows in this case. Similarly, another application of ˜4.7 in the case $b>0$ yields the existence of $\alpha>0$ such that $y^{b}(uu_{\nu}+vv_{\nu})\in C^{0,\alpha}(B_{r_{0}}^{+}\cup B_{r_{0}}^{\prime})$. The remainder of the argument follows as in the previous case. This completes the proof. ∎
@@
(5.12) $D_{0}^{\prime}(r)=\frac{n+b-1}{r}D_{0}(r)+2\int_{S_{r}^{+}}y^{b}(u_{\nu}^{2}+v_{\nu}^{2})\,d\mathcal{H}^{n}\\ -\frac{2}{r}\int_{B_{r}^{+}}y^{b}(X\cdot\nabla u)v\,dX-\frac{2}{r}\int_{B_{r}^{\prime}}(x\cdot\nabla_{x}^{\prime}v)f(x,u)\,dx$
@@
(5.14) $D_{0}^{\prime}(r)=\int_{S_{r}^{+}}y^{b}(|\nabla u|^{2}+|\nabla v|^{2})\,d\mathcal{H}^{n}$
@@
(5.16) $N_{0}^{\prime}(r)=\frac{D_{0}(r)}{H(r)}+\frac{rD_{0}^{\prime}(r)}{H(r)}-\frac{rD_{0}(r)H^{\prime}(r)}{H(r)^{2}}.$
@@
(5.17) $=-(n+b-1)\frac{D_{0}(r)}{H(r)}+\frac{rD_{0}^{\prime}(r)}{H(r)}-\frac{2rD(r)^{2}}{H(r)^{2}}+\frac{2rP(r)D(r)}{H(r)^{2}}.$
@@
(5.18) $N_{0}^{\prime}(r)\geq-\frac{2}{H(r)}\int_{B_{r}^{+}}y^{b}(X\cdot\nabla u)v\,dX-\frac{2}{H(r)}\int_{B_{r}^{\prime}}(x\cdot\nabla_{x}^{\prime}v)f(x,u)\,dx+\frac{2rP(r)D(r)}{H(r)^{2}}$
@@
(5.19) $\frac{2}{H(r)}\int_{B_{r}^{+}}y^{b}(X\cdot\nabla u)v\,dX\geq-C(N_{0}(r)+r^{2}).$
@@
(5.20) $\int_{B_{r}^{\prime}}(x\cdot\nabla_{x}^{\prime}v)f(x,u)\,dx=\frac{1}{r}\int_{S_{r}^{+}}(y(x\cdot\nabla_{x}^{\prime}v)-|x|^{2}v_{y})f(x,u)\,d\mathcal{H}^{n}\\ +\int_{B_{r}^{+}}[nf(x,u)v_{y}+(x\cdot\nabla_{x}^{\prime}f(x,u))v_{y}+(x\cdot\nabla_{x}^{\prime}u)\partial_{\zeta}f(x,u)v_{y}-(x\cdot\nabla_{x}^{\prime}v)\partial_{\zeta}f(x,u)u_{y}]\,dX.$
@@
(5.21) $\leq Cr\int_{S_{r}^{+}}y^{b}\left(r|\nabla v|^{2}+\frac{u^{2}}{r}\right)\,d\mathcal{H}^{n},$
@@
(5.22) $\left|\frac{1}{r}\int_{S_{r}^{+}}(y(x\cdot\nabla_{x}^{\prime}v)-|x|^{2}v_{y})f(x,u)\,d\mathcal{H}^{n}\right|\leq C(r^{2}D_{0}^{\prime}(r)+H(r)).$
@@
(5.23) $\left|\int_{B_{r}^{+}}nf(x,u)v_{y}\,dX\right|$ $\leq C(rD_{0}(r)+H(r)),$
(5.24) $\left|\int_{B_{r}^{+}}(x\cdot\nabla_{x}^{\prime}u)\partial_{\zeta}f(x,u)v_{y}-(x\cdot\nabla_{x}^{\prime}v)\partial_{\zeta}f(x,u)u_{y})\,dX\right|$ $\leq CrD_{0}(r),$
(5.25) $\left|\int_{B_{r}^{+}}(x\cdot\nabla_{x}^{\prime}f(x,u))v_{y}\,dX\right|$ $\leq Cr(D_{0}(r)+rH(r)).$
@@
(5.26) $\frac{2}{H(r)}\int_{B_{r}^{\prime}}(x\cdot\nabla_{x}^{\prime}v)f(x,u)\,dx\geq-C\left(\frac{r^{2}D_{0}^{\prime}(r)}{H(r)}+1+N_{0}(r)\right)$
@@
(5.28) $N_{0}^{\prime}(r)\geq-C\left(1+N_{0}(r)+N_{0}(r)^{2}+\frac{r^{2}D_{0}^{\prime}(r)}{H(r)}\right)$
@@
(5.33) $h(r)\coloneqq\frac{H(r)}{r^{n+b}}=\frac{1}{r^{n+b}}\int_{S_{r}^{+}}y^{b}(u^{2}+v^{2})\,d\mathcal{H}^{n}.$
@@
    where $C$ is a constant that depends on $\mu$ and on $\|u\|_{L^{\infty}(B_{r_{0}}^{\prime})}^{p-2}$ (when $p>2$). In particular, $h(r)=\mathcal{O}(r^{2\mu})$ as $r\to 0^{+}$.
@@
(5.34) $\frac{d}{dr}\log(r^{-2\mu}h(r))=\frac{2}{r}\left(N_{0}(r)-\mu+\frac{rP(r)}{H(r)}\right).$
@@
(5.36) $\geq\frac{2}{r}\left(-\frac{(1-Cr^{1-b})Cr(1+\mu)^{2}}{1+Cr(1+\mu)}-Cr^{1-b}(1+\mu)\right).$